### v2

# TENG++: Time-Evolving Natural Gradient for Solving PDEs With Deep Neural Nets under General Boundary Conditions

Abstract
Partial Differential Equations (PDEs) are central to modeling complex systems across physical, biological, and engineering domains, yet traditional numerical methods often struggle with high-dimensional or complex problems. Physics-Informed Neural Networks (PINNs) have emerged as an efficient alternative by embedding physics-based constraints into deep learning frameworks, but they face challenges in achieving high accuracy and handling complex boundary conditions. In this work, we extend the Time-Evolving Natural Gradient (TENG) framework to address Dirichlet boundary conditions, integrating natural gradient optimization with numerical time-stepping schemes, including Euler and Heun methods, to ensure both stability and accuracy. By incorporating boundary condition penalty terms into the loss function, the proposed approach enables precise enforcement of Dirichlet constraints. Experiments on the heat equation demonstrate the superior accuracy of the Heun method due to its second-order corrections and the computational efficiency of Euler’s method for simpler scenarios. This work establishes a foundation for extending the framework to Neumann and mixed boundary conditions, as well as broader classes of PDEs, advancing the applicability of neural network-based solvers for real-world problems.
Machine Learning, ICML
1 Introduction
Partial Differential Equations (PDEs) are foundational to modeling a wide range of physical, biological, and engineering systems. Solving PDEs is a cornerstone of computational science, but traditional numerical methods such as finite element and finite difference approaches can become computationally prohibitive, particularly in high-dimensional or complex domains. In recent years, Physics-Informed Neural Networks (PINNs)(Raissi:PINN) have emerged as a promising alternative, leveraging the expressive power of deep learning to approximate solutions to PDEs by embedding the governing equations into the network’s loss function.
Physics-Informed Neural Networks (PINNs) offer significant advantages, including being mesh-free and computationally efficient. These qualities make them particularly appealing for solving complex systems of equations like the Navier-Stokes equations, which have been a major focus of research in fluid dynamics. Despite significant progress over the past 50 years in the mathematical community, challenges remain: seamlessly incorporating noisy data into existing algorithms remains elusive, and generating meshes in high-dimensional spaces is inherently complex. In (Cai:mesh_free), the authors leveraged PINNs to address these challenges, demonstrating their effectiveness in solving a range of fluid dynamics problems. Remarkably, PINNs can also tackle ill-posed problems (Cai:ill_posed), which is typically infeasible using classical methods.
Despite their advantages of being mesh-free and computationally efficient, Physics-Informed Neural Networks (PINNs) have yet to achieve the level of accuracy offered by classical methods (Grossmann:accuracy; Cuomo:accuracy). This limitation is particularly evident in solving initial value problems in partial differential equations (PDEs), which are fundamental for modeling the evolution of dynamical systems. While machine learning techniques have shown promise in approximating solutions to complex partial differential equations (PDEs), they often struggle to achieve high accuracy, particularly when dealing with intricate initial conditions. This difficulty arises from the cumulative propagation of errors in PDE solvers over time, necessitating precise solutions at each time step to maintain overall accuracy.
Although various training strategies, such as global-in-time training (Muller2023) and sequential-in-time training (Chen2023a; Berman2023), have been developed to mitigate these challenges, achieving consistent accuracy remains a significant obstacle to the practical application of PINNs. Addressing this, the authors in (Chen:TENG) proposed the Time-Evolving Natural Gradient (TENG) method, which outperformed all previously benchmarked approaches, offering a promising advancement in the field.
In (Chen:TENG), the authors primarily focused on periodic boundary conditions. While effective in certain contexts, periodic boundary conditions are limited in scope. Non-periodic boundary conditions, such as Dirichlet or Neumann conditions, are essential for modeling a broader range of real-world phenomena where solutions are governed by fixed values or fluxes at the boundaries. These conditions enable accurate representation of complex systems, including heat transfer, fluid flow, and electromagnetic fields, where periodic assumptions are not applicable.
In this work, we extend the method proposed in (Chen:TENG) to handle arbitrary boundary conditions, broadening its applicability to a larger class of PDE problems. This enhancement significantly increases its potential for addressing real-world challenges. To validate the effectiveness of our approach, we present a series of numerical experiments on benchmark PDE problems of varying complexity.
2 Related Work
2.1 Machine Learning for Solving PDEs
Machine learning has emerged as a promising tool for solving partial differential equations (PDEs), utilizing neural networks as function approximators to model solutions. Broadly, there are two categories of approaches: global-in-time optimization and sequential-in-time optimization.
Global-in-time optimization, such as the Physics-Informed Neural Network (PINN) framework (Raissi2019), aims to optimize a neural network representation of the solution over the entire time and spatial domain simultaneously. Variants of this approach include the Deep Ritz Method (Weinan2017), which leverages the variational form of the PDE when it exists. While effective in many cases, these methods can struggle with maintaining accuracy, particularly in problems with intricate boundary conditions, where incorporating boundary constraints seamlessly remains challenging.
In contrast, sequential-in-time optimization, also known as the neural Galerkin method, focuses on updating the neural network representation step-by-step in time. Methods in this category include the Time-Dependent Variational Principle (TDVP) (Dirac1930; Koch2007; Carleo2017; Du2021; Berman2023) and Optimization-Based Time Integration (OBTI) (Kochkov2018; Gutierrez2022). Sequential-in-time strategies are particularly advantageous in handling evolving boundary conditions, as they allow for explicit incorporation of such conditions at each time step.
Beyond these direct PDE-solving strategies, data-driven approaches have also been employed to model PDEs based on observed data. These methods include Neural ODEs (Chen2018), Graph Neural Networks (Pfaff2020; SanchezGonzalez2020), Neural Fourier Operators (Li2020), and DeepONet (Lu2019). While these techniques have demonstrated strong performance in approximating PDE solutions, adapting them to accurately and efficiently enforce complex boundary conditions remains a significant challenge.
2.2 Natural Gradient Optimization
Natural gradient optimization, introduced by Amari (Amari1998), has become a cornerstone in second-order optimization techniques for machine learning. By adjusting the gradient direction based on the Fisher information matrix, natural gradient descent incorporates the geometry of the parameter space, leading to faster and more stable convergence compared to standard gradient methods. This feature makes it particularly effective for problems with complex constraints, such as boundary conditions in PDEs.
Natural gradient methods have been widely applied in neural network optimization (Pascanu2013), reinforcement learning (Kakade2001), quantum optimization, and PINN training (Muller2023). In the context of solving PDEs, these methods can improve the enforcement of boundary conditions by effectively navigating the optimization landscape shaped by the PDE constraints and boundary terms. Recent works have demonstrated the utility of natural gradients in PINN training (Muller2023), highlighting their potential to address issues such as cumulative error propagation and the incorporation of boundary constraints.
2.3 Boundary Conditions in PDEs and Neural Networks
Handling boundary conditions is a critical challenge in applying machine learning to PDEs. Global-in-time methods, such as PINNs, typically encode boundary conditions directly into the loss function. However, for complex or non-periodic boundary conditions, achieving accurate enforcement can become computationally challenging. Sequential-in-time methods, such as TDVP and OBTI, provide a more flexible framework for integrating boundary conditions, as they allow conditions to be explicitly updated at each time step. This stepwise approach is particularly suitable for problems involving dynamic or mixed boundary conditions, where the constraints may evolve over time.
Recent advancements in natural gradient optimization offer an additional layer of robustness in solving PDEs with boundary constraints. By leveraging second-order information, natural gradient methods can better balance the interplay between minimizing PDE residuals and satisfying boundary conditions, making them well-suited for addressing problems with intricate or non-periodic boundaries.
3 Methodology
The Time-Evolving Natural Gradient (TENG) framework (Chen:TENG) combines the time-dependent variational principle with optimization-based time integration using the natural gradient. This approach leverages second-order optimization to enhance convergence and numerical stability. While TENG has demonstrated effectiveness under periodic boundary conditions, it does not generalize to problems with arbitrary boundary conditions, such as Dirichlet or Neumann conditions. In this work, we extend the TENG framework to address these limitations.
3.1 Generalized Loss Functions
To handle Dirichlet and Neumann boundary conditions, we modify the TENG framework by incorporating penalty terms into the loss function. These terms enforce the boundary conditions in addition to minimizing the PDE residuals.
3.1.1 Dirichlet Boundary Conditions
For Dirichlet boundary conditions, where the solution is constrained to fixed values at the domain boundaries, the total loss function is defined as:
$\mathcal{L}_{\text{Dirichlet}}=\mathcal{L}_{\text{PDE}}+\lambda_{\text{Dirichlet}}\|u(x_{\text{boundary}})-u_{\text{Dirichlet}}\|^{2},$
where:
  • $\mathcal{L}_{\text{PDE}}$: The residual loss of the governing PDE, ensuring compliance with the physical laws.
  • $\|u(x_{\text{boundary}})-u_{\text{Dirichlet}}\|^{2}$: A penalty term enforcing the prescribed Dirichlet boundary values.
  • $\lambda_{\text{Dirichlet}}$: A weighting factor controlling the relative importance of the boundary condition term.
This formulation ensures that the solution satisfies both the PDE and the Dirichlet boundary conditions.
3.2 Optimization Process
The optimization process iteratively minimizes the total loss function for each time step of the PDE solution. This is achieved using the natural gradient, which projects the gradient of the loss function in the $u$-space onto the parameter $\theta$-space. By formulating the parameter updates as a least-squares problem, the natural gradient method achieves second-order optimization, improving both stability and convergence rates.
The details of the parameter update process are implemented in the TENG_stepper subroutine, described in Algorithm 1.
Algorithm 1 TENG_stepper: Parameter Update Subroutine
Input: Initial parameters $\theta_{\text{init}}$, target solution $u_{\text{target}}$
 Initialize $n\leftarrow 0$, $\theta\leftarrow\theta_{\text{init}}$
repeat
  Compute the gradient: $\Delta u(x)\leftarrow-\alpha_{n}\frac{\partial L(\hat{u}_{\theta},u_{\text{target}})}{\partial\hat{u}_{\theta}}$
  Compute the Jacobian: $J(x)_{j}\leftarrow\frac{\partial\hat{u}_{\theta}(x)}{\partial\theta_{j}}$
  Solve the least-squares problem:
$\Delta\theta=\arg\min_{\Delta\theta\in\mathbb{R}^{N_{p}}}\left\|\Delta u(\cdot)-\sum_{j}J(\cdot)_{j}\Delta\theta_{j}\right\|^{2}_{L^{2}(X)}$
  Update parameters: $\theta\leftarrow\theta+\Delta\theta$
  Increment iteration: $n\leftarrow n+1$
until $n\geq N_{\text{it}}$
Output: $\theta$
3.3 Time-Stepping Schemes
The TENG_stepper subroutine is integrated into two numerical time-stepping schemes: Euler and Heun methods. These schemes iteratively evolve the parameters $\theta$ at each time step while ensuring numerical stability and accuracy.
3.3.1 Euler Method
Euler’s method is a first-order integration scheme that updates the solution at each time step based on the PDE residual. The algorithm is summarized in Algorithm 2.
Algorithm 2 TENG_Euler: A 1st-Order Integration Scheme
Input: Initial parameters $\theta_{t=0}$, time step $\Delta t$, total time $T$
 Initialize $t\leftarrow 0$
repeat
  Compute $u_{\text{target}}(x)\leftarrow\hat{u}_{\theta_{t}}(x)+\Delta tL\hat{u}_{\theta_{t}}(x)$
  Update $\theta_{t+\Delta t}\leftarrow\text{TENG\_stepper}(\theta_{t},u_{\text{target}})$
  Increment $t\leftarrow t+\Delta t$
until $t\geq T$
Output: Final parameters $\theta_{t=T}$
3.3.2 Heun Method
Heun’s method is a second-order integration scheme that incorporates an intermediate step to improve accuracy. The algorithm is summarized in Algorithm 3.
Algorithm 3 TENG_Heun: A 2nd-Order Integration Scheme
Input: Initial parameters $\theta_{t=0}$, time step $\Delta t$, total time $T$
 Initialize $t\leftarrow 0$
repeat
  Compute $u_{\text{temp}}(x)\leftarrow\hat{u}_{\theta_{t}}(x)+\Delta tL\hat{u}_{\theta_{t}}(x)$
  Update $\theta_{\text{temp}}\leftarrow\text{TENG\_stepper}(\theta_{t},u_{\text{temp}})$
  Compute $u_{\text{target}}(x)\leftarrow\hat{u}_{\theta_{t}}(x)+\frac{\Delta t}{2}\big(L\hat{u}_{\theta_{t}}(x)+L\hat{u}_{\theta_{\text{temp}}}(x)\big)$
  Update $\theta_{t+\Delta t}\leftarrow\text{TENG\_stepper}(\theta_{\text{temp}},u_{\text{target}})$
  Increment $t\leftarrow t+\Delta t$
until $t\geq T$
Output: Final parameters $\theta_{t=T}$
3.4 Numerical Experiments
We validated the proposed framework by solving the heat equation with Dirichlet boundary conditions. The experiments compared the Euler and Heun methods in terms of accuracy, stability, and convergence. The setup included:
  1. 1.
    Solving the heat equation on a circular domain with analytical solutions derived from Bessel functions.
  2. 2.
    Testing various initial conditions to evaluate robustness and precision.
The results demonstrated that the Heun method achieves higher accuracy due to its second-order corrections, while the Euler method provides a computationally efficient alternative for simpler scenarios.
4 Experiments
We tested our algorithm on the heat equation with Dirichlet boundary conditions. The heat equation is a fundamental partial differential equation that models the distribution of heat (or diffusion of substances) in a given region over time, making it essential in fields like physics, engineering, and materials science. Its mathematical framework also provides key insights into broader areas such as probability theory, financial modeling, and the study of partial differential equations in general. Specifically, let $\Omega=B(0,1)\subset\mathbb{R}^{2}$ be the open disk centered on 0 with radius 1. We consider the two-dimensional isotropic heat equation
$\frac{\partial u}{\partial t}=\nu(\frac{\partial^{2}u}{\partial x_{1}^{2}}+\frac{\partial^{2}u}{\partial x_{2}^{2}})$
with a diffusivity constant $\nu=1/10$. We also impose the Dirichlet boundary condition that
$u(x,t)=0,\forall x\in\partial\Omega.$
Since the heat equation on a disk has an analytical solution with Bessel functions as the basis, we consider initial conditions as linear combinations of Bessel functions. This approach provides precise analytical solutions, making it ideal for benchmarking.
4.1 Experiment 1
We consider the initial condition
$u_{0}(r,\theta)=\frac{1}{4}(Z_{01}(r,\theta)-\frac{1}{4}Z_{02}(r,\theta)+\frac{1}{16}Z_{03}(r,\theta)\\ -\frac{1}{64}Z_{04}(r,\theta)+Z_{11}(r,\theta)-\frac{1}{2}Z_{12}(r,\theta)+\frac{1}{4}Z_{13}(r,\theta)\\ -\frac{1}{8}Z_{14}(r,\theta)+Z_{21}(r,\theta)+Z_{31}(r,\theta)+Z_{41}(r,\theta))$ (1)
where $r$ and $\theta$ are the polar coordinate variables and $Z_{mn}$ represent the disk harmonics defined as $Z_{mn}(r,\theta)=J_{m}(\lambda_{nm}r)\cos(m\theta)$ with $J_{m}$ the $m$th Bessel function and $\lambda_{mn}$ the $n$th zero of the $m$th Bessel function.
We solve this equation on the domain $B(0,1)\times[0,\mathcal{T}]$, where $\mathcal{T}=4$, using TENG_Heun methods with a stepsize of $h=0.005$. We also output the result on the same domain with $\mathcal{T}=0.8$ using TENG_Euler with the same stepsize $h=0.005$. The resulting errors of TENG_Heun and TENG_Euler are visualized in Figure 1 and Figure 2 respectively.
Figure 1: TENG_Heun Result
Figure 2: TENG_Euler Result
To initialize the neural network, we utilized a set of supervised pre-trained weights. Our results show that the error for TENG Euler is significantly larger compared to TENG Heun. Notably, TENG Heun demonstrates promising accuracy, with error already within 1e41e-4. We expect more improvement is possible if using a higher-order integration scheme, such as RK4.
4.2 Experiment 2
We consider the same initial condition
$u_{0}(r,\theta)=\frac{1}{4}(Z_{01}(r,\theta)-\frac{1}{4}Z_{02}(r,\theta)+\frac{1}{16}Z_{03}(r,\theta)\\ -\frac{1}{64}Z_{04}(r,\theta)+Z_{11}(r,\theta)-\frac{1}{2}Z_{12}(r,\theta)+\frac{1}{4}Z_{13}(r,\theta)\\ -\frac{1}{8}Z_{14}(r,\theta)+Z_{21}(r,\theta)+Z_{31}(r,\theta)+Z_{41}(r,\theta))$ (2)
Once again, we solve this equation on the domain $B(0,1)\times[0,\mathcal{T}]$, where $\mathcal{T}=4$ using the TENG_Huen method and $\mathcal{T}=0.8$ using the TENG_Euler method. However, instead of importing a set of pre-trained weights for neural network initialization, we designed the model to consist of two neural networks. The model’s output is computed as the difference between the two neural networks, plus the initial condition.
During initialization, both neural networks were set to have equal weights. During training, the parameters of one neural network were frozen, and only the parameters of the other network were updated. This approach resulted in significantly larger errors compared to the pre-trained initialization.
These results highlight the critical importance of the initialization phase in guiding the model towards the correct solution. The choice of initial parameters plays a pivotal role in ensuring effective training and reducing error.
5 Discussion
In this work, we generalized periodic boundary conditions in (Chen:TENG) to Dirichlet boundary conditions, demonstrating the adaptability of our approach to handle different types of constraints. This extension broadens the applicability of the method to a wider range of physical and mathematical problems, where Dirichlet boundary conditions play a critical role. The results underscore the potential of this generalized framework for accurately solving boundary value problems while maintaining computational efficiency and flexibility.
Looking ahead, several exciting directions emerge for future research. An immediate extension would be the incorporation of Neumann boundary conditions, which model flux-constrained scenarios and are highly relevant in heat transfer, fluid dynamics, and electromagnetics. Exploring mixed boundary conditions, where different segments of the boundary enforce Dirichlet or Neumann constraints, would further enhance the utility of the approach for real-world applications.
Additionally, extending the current schema to a broader class of partial differential equations (PDEs) beyond initial-value problems would be an important advancement. Addressing PDEs that involve complex geometries, nonlinearity, or time-dependent boundaries could significantly expand the scope of the method. Such generalizations would not only enhance the mathematical framework but also provide critical tools for scientific and engineering applications, ranging from materials science to climate modeling.
These advancements would represent significant contributions to the field, offering both theoretical insights and practical tools for solving challenging problems across disciplines.
6 Acknowledgments
We would like to thank Professor Di Luo and Zhuo Chen, their valuable suggestions and great support throughout this project.
Appendix A Hyperparamter
Table 1: Hyperparameters for Euler and Heun Methods
Hyperparameter Euler Method Heun Method
D 0.1 0.1
Equation Heat Heat
Number of Steps 800 800
Iterations per Step 5 5
Time Step (dt) 0.001 0.005
Integrator Euler Heun
Save Directory None None
Model State model_state_circ_2d.pickle model_state_circ_2d.pickle
Model Seed 1234 1234
Number of Samples 65536 65536
Sampler Seed 4321 4321
Policy Gradient Params 1536 1536
Policy Gradient Seed 8844 8844
Policy Gradient 2 Params 1024 1024
Policy Gradient 2 Seed 8848 8848

### v3

@@
-TENG++: Time-Evolving Natural Gradient for Solving PDEs With Deep Neural Nets under General Boundary Conditions
+Solving PDEs With Deep Neural Nets under General Boundary Conditions
 Abstract
@@
 This work establishes a foundation for extending the framework to Neumann and mixed boundary conditions, as well as broader classes of PDEs, advancing the applicability of neural network-based solvers for real-world problems.
-Machine Learning, ICML
 1 Introduction
@@
 Solving PDEs is a cornerstone of computational science, but traditional numerical methods such as finite element and finite difference approaches can become computationally prohibitive, particularly in high-dimensional or complex domains.
-In recent years, Physics-Informed Neural Networks (PINNs)(Raissi:PINN) have emerged as a promising alternative, leveraging the expressive power of deep learning to approximate solutions to PDEs by embedding the governing equations into the network’s loss function.
+In recent years, Physics-Informed Neural Networks (PINNs)Raissi et al.
+(2017) have emerged as a promising alternative, leveraging the expressive power of deep learning to approximate solutions to PDEs by embedding the governing equations into the network’s loss function.
 Physics-Informed Neural Networks (PINNs) offer significant advantages, including being mesh-free and computationally efficient.
@@
 Despite significant progress over the past 50 years in the mathematical community, challenges remain: seamlessly incorporating noisy data into existing algorithms remains elusive, and generating meshes in high-dimensional spaces is inherently complex.
-In (Cai:mesh_free), the authors leveraged PINNs to address these challenges, demonstrating their effectiveness in solving a range of fluid dynamics problems.
-Remarkably, PINNs can also tackle ill-posed problems (Cai:ill_posed), which is typically infeasible using classical methods.
-Despite their advantages of being mesh-free and computationally efficient, Physics-Informed Neural Networks (PINNs) have yet to achieve the level of accuracy offered by classical methods (Grossmann:accuracy; Cuomo:accuracy).
+In Cai et al.
+(2021a), the authors leveraged PINNs to address these challenges, demonstrating their effectiveness in solving a range of fluid dynamics problems.
+Remarkably, PINNs can also tackle ill-posed problems Cai et al.
+(2021b), which is typically infeasible using classical methods.
+Despite their advantages of being mesh-free and computationally efficient, Physics-Informed Neural Networks (PINNs) have yet to achieve the level of accuracy offered by classical methods Grossmann et al.
+(2023); Cuomo et al.
+(2022).
 This limitation is particularly evident in solving initial value problems in partial differential equations (PDEs), which are fundamental for modeling the evolution of dynamical systems.
@@
 This difficulty arises from the cumulative propagation of errors in PDE solvers over time, necessitating precise solutions at each time step to maintain overall accuracy.
-Although various training strategies, such as global-in-time training (Muller2023) and sequential-in-time training (Chen2023a; Berman2023), have been developed to mitigate these challenges, achieving consistent accuracy remains a significant obstacle to the practical application of PINNs.
-Addressing this, the authors in (Chen:TENG) proposed the Time-Evolving Natural Gradient (TENG) method, which outperformed all previously benchmarked approaches, offering a promising advancement in the field.
-In (Chen:TENG), the authors primarily focused on periodic boundary conditions.
+Although various training strategies, such as global-in-time training Müller and Zeinhofer (2023) and sequential-in-time training Chen et al.
+(2023); Berman and Peherstorfer (2023), have been developed to mitigate these challenges, achieving consistent accuracy remains a significant obstacle to the practical application of PINNs.
+Addressing this, the authors in Chen et al.
+(2024) proposed the Time-Evolving Natural Gradient (TENG) method, which outperformed all previously benchmarked approaches, offering a promising advancement in the field.
+In Chen et al.
+(2024), the authors primarily focused on periodic boundary conditions.
 While effective in certain contexts, periodic boundary conditions are limited in scope.
@@
 These conditions enable accurate representation of complex systems, including heat transfer, fluid flow, and electromagnetic fields, where periodic assumptions are not applicable.
-In this work, we extend the method proposed in (Chen:TENG) to handle arbitrary boundary conditions, broadening its applicability to a larger class of PDE problems.
+In this work, we extend the method proposed in Chen et al.
+(2024) to handle arbitrary boundary conditions, broadening its applicability to a larger class of PDE problems.
 This enhancement significantly increases its potential for addressing real-world challenges.
@@
 Broadly, there are two categories of approaches: global-in-time optimization and sequential-in-time optimization.
-Global-in-time optimization, such as the Physics-Informed Neural Network (PINN) framework (Raissi2019), aims to optimize a neural network representation of the solution over the entire time and spatial domain simultaneously.
-Variants of this approach include the Deep Ritz Method (Weinan2017), which leverages the variational form of the PDE when it exists.
+Global-in-time optimization, such as the Physics-Informed Neural Network (PINN) framework Raissi et al.
+(2019), aims to optimize a neural network representation of the solution over the entire time and spatial domain simultaneously.
+Variants of this approach include the Deep Ritz Method E and Yu (2017), which leverages the variational form of the PDE when it exists.
 While effective in many cases, these methods can struggle with maintaining accuracy, particularly in problems with intricate boundary conditions, where incorporating boundary constraints seamlessly remains challenging.
 In contrast, sequential-in-time optimization, also known as the neural Galerkin method, focuses on updating the neural network representation step-by-step in time.
-Methods in this category include the Time-Dependent Variational Principle (TDVP) (Dirac1930; Koch2007; Carleo2017; Du2021; Berman2023) and Optimization-Based Time Integration (OBTI) (Kochkov2018; Gutierrez2022).
+Methods in this category include the Time-Dependent Variational Principle (TDVP) Dirac (1930); Koch and Lubich (2007); Carleo and Troyer (2017); Du and Zaki (2021); Berman and Peherstorfer (2023) and Optimization-Based Time Integration (OBTI) Kochkov and Clark (2018); Gutierrez and Mendl (2022).
 Sequential-in-time strategies are particularly advantageous in handling evolving boundary conditions, as they allow for explicit incorporation of such conditions at each time step.
 Beyond these direct PDE-solving strategies, data-driven approaches have also been employed to model PDEs based on observed data.
-These methods include Neural ODEs (Chen2018), Graph Neural Networks (Pfaff2020; SanchezGonzalez2020), Neural Fourier Operators (Li2020), and DeepONet (Lu2019).
+These methods include Neural ODEs Chen et al.
+(2018), Graph Neural Networks Pfaff et al.
+(2020); Sanchez-Gonzalez et al.
+(2020), Neural Fourier Operators Li et al.
+(2020), and DeepONet Lu et al.
+(2019).
 While these techniques have demonstrated strong performance in approximating PDE solutions, adapting them to accurately and efficiently enforce complex boundary conditions remains a significant challenge.
 2.2 Natural Gradient Optimization
-Natural gradient optimization, introduced by Amari (Amari1998), has become a cornerstone in second-order optimization techniques for machine learning.
+Natural gradient optimization, introduced by Amari Amari (1998), has become a cornerstone in second-order optimization techniques for machine learning.
 By adjusting the gradient direction based on the Fisher information matrix, natural gradient descent incorporates the geometry of the parameter space, leading to faster and more stable convergence compared to standard gradient methods.
 This feature makes it particularly effective for problems with complex constraints, such as boundary conditions in PDEs.
-Natural gradient methods have been widely applied in neural network optimization (Pascanu2013), reinforcement learning (Kakade2001), quantum optimization, and PINN training (Muller2023).
+Natural gradient methods have been widely applied in neural network optimization Pascanu et al.
+(2013), reinforcement learning Kakade (2001), quantum optimization, and PINN training Müller and Zeinhofer (2023).
 In the context of solving PDEs, these methods can improve the enforcement of boundary conditions by effectively navigating the optimization landscape shaped by the PDE constraints and boundary terms.
-Recent works have demonstrated the utility of natural gradients in PINN training (Muller2023), highlighting their potential to address issues such as cumulative error propagation and the incorporation of boundary constraints.
+Recent works have demonstrated the utility of natural gradients in PINN training Müller and Zeinhofer (2023), highlighting their potential to address issues such as cumulative error propagation and the incorporation of boundary constraints.
 2.3 Boundary Conditions in PDEs and Neural Networks
@@
 3 Methodology
-The Time-Evolving Natural Gradient (TENG) framework (Chen:TENG) combines the time-dependent variational principle with optimization-based time integration using the natural gradient.
+The Time-Evolving Natural Gradient (TENG) framework Chen et al.
+(2024) combines the time-dependent variational principle with optimization-based time integration using the natural gradient.
 This approach leverages second-order optimization to enhance convergence and numerical stability.
@@
 By formulating the parameter updates as a least-squares problem, the natural gradient method achieves second-order optimization, improving both stability and convergence rates.
-The details of the parameter update process are implemented in the TENG_stepper subroutine, described in Algorithm 1.
+The details of the parameter update process are implemented in the TENG_stepper subroutine, described in 1.
 Algorithm 1 TENG_stepper: Parameter Update Subroutine
@@
 repeat
-  Compute the gradient: $\Delta u(x)\leftarrow-\alpha_{n}\frac{\partial L(\hat{u}_{\theta},u_{\text{target}})}{\partial\hat{u}_{\theta}}$
-  Compute the Jacobian: $J(x)_{j}\leftarrow\frac{\partial\hat{u}_{\theta}(x)}{\partial\theta_{j}}$
+  Compute the gradient: $\Delta u(x)\leftarrow-\alpha_{n}\frac{\partial L(\widehat{u}_{\theta},u_{\text{target}})}{\partial\widehat{u}_{\theta}}$
+  Compute the Jacobian: $J(x)_{j}\leftarrow\frac{\partial\widehat{u}_{\theta}(x)}{\partial\theta_{j}}$
   Solve the least-squares problem:
@@
 3.3.1 Euler Method
-Euler’s method is a first-order integration scheme that updates the solution at each time step based on the PDE residual. The algorithm is summarized in Algorithm 2.
+Euler’s method is a first-order integration scheme that updates the solution at each time step based on the PDE residual. The algorithm is summarized in 2.
 Algorithm 2 TENG_Euler: A 1st-Order Integration Scheme
@@
 repeat
-  Compute $u_{\text{target}}(x)\leftarrow\hat{u}_{\theta_{t}}(x)+\Delta tL\hat{u}_{\theta_{t}}(x)$
+  Compute $u_{\text{target}}(x)\leftarrow\widehat{u}_{\theta_{t}}(x)+\Delta tL\widehat{u}_{\theta_{t}}(x)$
   Update $\theta_{t+\Delta t}\leftarrow\text{TENG\_stepper}(\theta_{t},u_{\text{target}})$
@@
 3.3.2 Heun Method
-Heun’s method is a second-order integration scheme that incorporates an intermediate step to improve accuracy. The algorithm is summarized in Algorithm 3.
+Heun’s method is a second-order integration scheme that incorporates an intermediate step to improve accuracy. The algorithm is summarized in 3.
 Algorithm 3 TENG_Heun: A 2nd-Order Integration Scheme
@@
 repeat
-  Compute $u_{\text{temp}}(x)\leftarrow\hat{u}_{\theta_{t}}(x)+\Delta tL\hat{u}_{\theta_{t}}(x)$
+  Compute $u_{\text{temp}}(x)\leftarrow\widehat{u}_{\theta_{t}}(x)+\Delta tL\widehat{u}_{\theta_{t}}(x)$
   Update $\theta_{\text{temp}}\leftarrow\text{TENG\_stepper}(\theta_{t},u_{\text{temp}})$
-  Compute $u_{\text{target}}(x)\leftarrow\hat{u}_{\theta_{t}}(x)+\frac{\Delta t}{2}\big(L\hat{u}_{\theta_{t}}(x)+L\hat{u}_{\theta_{\text{temp}}}(x)\big)$
+  Compute $u_{\text{target}}(x)\leftarrow\widehat{u}_{\theta_{t}}(x)+\frac{\Delta t}{2}\big(L\widehat{u}_{\theta_{t}}(x)+L\widehat{u}_{\theta_{\text{temp}}}(x)\big)$
   Update $\theta_{t+\Delta t}\leftarrow\text{TENG\_stepper}(\theta_{\text{temp}},u_{\text{target}})$
@@
 We consider the initial condition
-$u_{0}(r,\theta)=\frac{1}{4}(Z_{01}(r,\theta)-\frac{1}{4}Z_{02}(r,\theta)+\frac{1}{16}Z_{03}(r,\theta)\\ -\frac{1}{64}Z_{04}(r,\theta)+Z_{11}(r,\theta)-\frac{1}{2}Z_{12}(r,\theta)+\frac{1}{4}Z_{13}(r,\theta)\\ -\frac{1}{8}Z_{14}(r,\theta)+Z_{21}(r,\theta)+Z_{31}(r,\theta)+Z_{41}(r,\theta))$ (1)
+$u_{0}(r,\theta)$ $=\frac{1}{4}(Z_{01}(r,\theta)-\frac{1}{4}Z_{02}(r,\theta)+\frac{1}{16}Z_{03}(r,\theta)$
+$-\frac{1}{64}Z_{04}(r,\theta)+Z_{11}(r,\theta)-\frac{1}{2}Z_{12}(r,\theta)+\frac{1}{4}Z_{13}(r,\theta)$
+$-\frac{1}{8}Z_{14}(r,\theta)+Z_{21}(r,\theta)+Z_{31}(r,\theta)+Z_{41}(r,\theta))$
 where $r$ and $\theta$ are the polar coordinate variables and $Z_{mn}$ represent the disk harmonics defined as $Z_{mn}(r,\theta)=J_{m}(\lambda_{nm}r)\cos(m\theta)$ with $J_{m}$ the $m$th Bessel function and $\lambda_{mn}$ the $n$th zero of the $m$th Bessel function.
@@
 We consider the same initial condition
-$u_{0}(r,\theta)=\frac{1}{4}(Z_{01}(r,\theta)-\frac{1}{4}Z_{02}(r,\theta)+\frac{1}{16}Z_{03}(r,\theta)\\ -\frac{1}{64}Z_{04}(r,\theta)+Z_{11}(r,\theta)-\frac{1}{2}Z_{12}(r,\theta)+\frac{1}{4}Z_{13}(r,\theta)\\ -\frac{1}{8}Z_{14}(r,\theta)+Z_{21}(r,\theta)+Z_{31}(r,\theta)+Z_{41}(r,\theta))$ (2)
+$u_{0}(r,\theta)$ $=\frac{1}{4}(Z_{01}(r,\theta)-\frac{1}{4}Z_{02}(r,\theta)+\frac{1}{16}Z_{03}(r,\theta)-\frac{1}{64}Z_{04}(r,\theta)$
+$+Z_{11}(r,\theta)-\frac{1}{2}Z_{12}(r,\theta)+\frac{1}{4}Z_{13}(r,\theta)$
+$-\frac{1}{8}Z_{14}(r,\theta)+Z_{21}(r,\theta)+Z_{31}(r,\theta)+Z_{41}(r,\theta))$
 Once again, we solve this equation on the domain $B(0,1)\times[0,\mathcal{T}]$, where $\mathcal{T}=4$ using the TENG_Huen method and $\mathcal{T}=0.8$ using the TENG_Euler method.
@@
 5 Discussion
-In this work, we generalized periodic boundary conditions in (Chen:TENG) to Dirichlet boundary conditions, demonstrating the adaptability of our approach to handle different types of constraints.
+In this work, we generalized periodic boundary conditions in Chen et al.
+(2024) to Dirichlet boundary conditions, demonstrating the adaptability of our approach to handle different types of constraints.
 This extension broadens the applicability of the method to a wider range of physical and mathematical problems, where Dirichlet boundary conditions play a critical role.
@@
 6 Acknowledgments
-We would like to thank Professor Di Luo and Zhuo Chen, their valuable suggestions and great support throughout this project.
+I would like to thank Professor Di Luo and Zhuo Chen, their valuable suggestions and great support throughout this project.
+This note was not reviewed, approved, or endorsed by Prof.
+Di Luo or Zhuo Chen.
+Any errors are solely the author’s responsibility.
+I would like to thank Xinjie He for her valuable contribution.
 Appendix A Hyperparamter